OpenStudy (anonymous):

PLEASE HELP Find the number of real solutions to the functions below by graphing. 7) x + 3 = (x – 5)^2 8) x4 = x – 1 9) x3 = 5 – x

OpenStudy (danjs):

the first one you can maybe think of it as intersection points for the line x + 3 and the parabola (x-5)^2

OpenStudy (danjs):

or you can graph the parabola, and show where it crosses the x axis

OpenStudy (danjs):

ax^2+bx+c = 0

OpenStudy (danjs):

7) goes to x^2 -11x +22 =0 ,

OpenStudy (anonymous):

im allowed on use the sight https://www.desmos.com/calculator to graph it will you help me so i know what to type in each box im sorry im horrible at math lol

OpenStudy (danjs):

graph that parabola y = f(x) and show the x values where y = f(x) = 0

OpenStudy (danjs):

k, for 7) graph y=x^2-11x+22 show the x values when y is zero (x-axis) those are the solutions

OpenStudy (danjs):

OR Graph y=x+3 and y = (x-5)^2 and they should cross two times and have intersection points, the x values are the solutions

OpenStudy (anonymous):

okay i graphed 7 thank you!

OpenStudy (anonymous):

so how would i do numbers 8 and 9 now? im just confused on how to rearrange them like you did

OpenStudy (danjs):

show where x^4 - x + 1 crosses the x axis

OpenStudy (danjs):

y = x^4 - x + 1 if it even does

OpenStudy (triciaal):

another approach do a table with x values (common -3 to 3) calculate the corresponding y value plot the points "see" the intersection

OpenStudy (danjs):

i think that is what they want, as @triciaal says,, just show how many times y = f(x) froses the x axis, number of roots, just "see" it

OpenStudy (danjs):

the point to know the general shape of each of those degrees

OpenStudy (anonymous):

okay i graphed number 8 its correct thank you

OpenStudy (danjs):

no solution

OpenStudy (anonymous):

what would number 7 be?

OpenStudy (triciaal):